OpenStudy (anonymous):

HELP? pleaseeee. use the quotient rule to differentiate the function….. f(t)=cost/t^3

OpenStudy (anonymous):

Easy it's (cost) / t^3 right? (cost)(t^-3) = ((-sint)(t^-3) + (cost)(-3t^-4)) / (t^-3)^2 = ((-sint)(t^-3) + (cost)(-3t^-4)) / (t^-6)

OpenStudy (anonymous):

yessss, why -3 though?

OpenStudy (cwrw238):

yes its [ (-sin t) * t^3 - cos t * 3t^2 ] / t^6

OpenStudy (anonymous):

ohhh ok ok

OpenStudy (anonymous):

because the power rule brings t^-3 to -3t^-4. t^-3 = 1/t^3

OpenStudy (anonymous):

I also nee help with this one…. - find f'(x) and f'(c ) f(x)=(x^2-2x+1)(x^3-1) c=1

OpenStudy (cwrw238):

u can simplify t^2[- t sint - 3 cos t] / t^6 = [-t sint - 3cost] / t^4

OpenStudy (cwrw238):

f(x)=(x^2-2x+1)(x^3-1) use the product rule f'(x) = 3x^2(x^2 - 2x + 1) + (x^3 - 1)(2x - 2) f'(c) = 3(1 - 2 + 1) + (1 - 1)(2 - 2) when c = 1 = 3(0) + 0 = 0

OpenStudy (anonymous):

thankyou so much. I actually get how you got that now:) thanks!

OpenStudy (anonymous):

@cwrw238